# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

### v2

@@ III.4 Type-II conserved currents and non-invertible symmetry operators
Besides the type-I equations of motion discussed above, the theory also admits another class of EoMs. As an example, consider a local variation of $a^{1}$, $a^{1}\rightarrow a^{1}+\delta a^{1}$. The variation of the action is $\delta S\left[\delta a^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta a^{1}db^{1}+q\delta a^{1}a^{2}b^{3}$. Requiring $\delta S\left[\delta a^{1}\right]=0\mod{2\pi}$ yields
@@
A natural question is whether one can still extract a continuity equation from the type-II EoM (26). The answer is affirmative, but only under additional conditions. The key step is to rewrite the EoM into the form of a vanishing total derivative. To this end, we introduce two auxiliary fields $\phi^{2}$ and $\beta^{3}$ such that $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. We notice that $d\phi^{2}$ ($d\beta^{3}$) is an integral multiple of $a^{2}$ ($b^{3}$). Therefore, the fields $\phi^{2}$ and $\beta^{3}$ satisfy the Dirac quantization condition $\oint d\phi^{2}\in 2\pi\mathbb{Z}$ and $\oint d\beta^{3}\in 2\pi\mathbb{Z}$ when $\oint a^{2}$ and $\oint b^{3}$ take specific values. With $\phi^{2}$ and $\beta^{3}$, the EoM (26) can be rewritten as
@@
Here, the two coefficients $\frac{1}{2}$ are not the result of the gauge invariance: the above equation still holds if one chooses two coefficient $c_{1}$ and $c_{2}$ with $c_{1}+c_{2}=1$. Instead, we split the term $\frac{2\pi qa^{2}b^{3}}{N_{1}}$ in a symmetric way in order to respect the fact that the twisted term in the action (14) is symmetric under exchanging $a^{1}$ and $a^{2}$, up to an overall anti-symmetry of the wedge product. In this way, the currents extracted from the EoMs have the same symmetric formalism, see Table 1. Using $\frac{N_{2}}{2\pi}da^{2}=0$ and $\frac{N_{3}}{2\pi}db^{3}=0$, we further obtain
@@
$d\left(*J_{m}^{1}\right)=0,\ J_{m}^{1}=*\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right),$ (29)
@@
However, an important subtlety arises: while the EoM (26) holds identically, the continuity equation (29)—and thus the conservation of $J^{1}_{m}$—is valid only when $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$ are satisfied. Therefore, for $Q_{m}^{1}$ to be conserved, these relations must hold at least on the submanifold $\sigma$. Equivalently, these conditions can be expressed as $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$, because on the closed manifold $\sigma$ one has $\int_{\sigma}d\beta^{3}=0$, and $\int_{\gamma}d\phi^{2}=0$ for any closed curve $\gamma\subset\sigma$. In other words, although $\mathcal{U}_{1}$ is a unitary operator originating from $Q_{m}^{1}$, it defines a symmetry operation only when $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$ are satisfied. Accordingly, we define the symmetry operator as
@@
We further note that $L^{1}\left(\sigma\right)$ closely resembles the Wilson operator describing a loop excitation carrying one unit of $\mathbb{Z}_{N_{1}}$ gauge flux [76], up to an overall normalization factor. With this Wilson-operator picture in mind, the symmetry operator admits the following interpretation. Inserting such an operator creates a pair of loop excitations, transports one of them along $\sigma$, and finally annihilates the pair. During this process, the operator measures how many $\mathbb{Z}_{N_{1}}$ gauge charges (carried by particle excitations) are enclosed in the volume bounded by $\sigma$. This is analogous to the 22d toric code, where moving an $m$ excitation along a closed path measures the number of enclosed $e$ excitations. The enclosed number of $\mathbb{Z}_{N_{1}}$ gauge charges, $n_{e_{1}}$, is reflected by the phase factor $\exp\left({\rm i}\frac{2\pi}{N_{1}}\times n_{e_{1}}\right)$. Since the loop carries one unit of gauge flux, the measurement outcome is defined modulo $N_{1}$. The “symmetry” here is reflected in the fact that $\sigma$ can be smoothly deformed without changing $Q_{m}^{1}$, but only within the subspace where $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$. Outside this subspace, the symmetry need not exist. In this way, we establish a direct connection between a non-Abelian loop excitation and a non-invertible symmetry.
@@
$\mathcal{U}_{m_{1}}=e^{{\rm i}Q^{1}_{m_{1}}}=e^{{\rm i}\left(\int_{\sigma}m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (36)
@@
In the same manner, we can obtain the EoM of gauge field $b^{2}$,
@@
$d\left(*J_{m}^{2}\right)=0,\ J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (38)
@@
$L_{m_{2}}^{2}\left(\sigma\right)=\mathcal{U}_{m_{2}}\delta\left(\int_{\sigma}\frac{m_{2}pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{2}pN_{2}}{N_{123}}a^{2}\right),$ (41)
@@
$\mathcal{U}_{m_{2}}=$ $e^{{\rm i}\int_{\sigma}*J_{m_{2}}^{2}}=e^{{\rm i}\left(\int_{\sigma}m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right)}.$ (42)
@@
The above discussion also applies to the EoM of gauge field $a^{3}$,
@@
$d\left(*J_{m}^{3}\right)=0,\ J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right),$ (44)
@@
$\mathcal{U}_{m_{3}}=$ $e^{{\rm i}\int_{\gamma}*J_{m_{3}}^{3}}=e^{{\rm i}\left(\int_{\gamma}m_{3}a^{2}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right)}.$ (48)
@@ IV Fusion rules of symmetry operators
In this section, we investigate the fusion rules of the symmetry operators identified in Sec. III. Fusion rules characterize the outcome of consecutively applying two symmetry operators supported on the same submanifold, and they encode the underlying algebraic structure of a generalized symmetry. From a field-theoretical perspective, fusion is realized by inserting symmetry operators sequentially into correlation functions.
For invertible symmetries, fusion rules are group-like: each symmetry operator is labeled by a group element, and their fusion follows the group multiplication law. By contrast, for non-invertible symmetries, fusion rules are intrinsically non-group-like. In particular, the fusion of two non-invertible symmetry operators can produce a linear combination of several symmetry operators, rather than a single one.
To analyze fusion rules in a systematic manner, it is crucial to first identify the complete set of symmetry operators in the theory. In Sec. IV.1, we show that additional symmetry operators can be generated from new conserved currents, which themselves arise from linear combinations of existing conserved currents. This construction elucidates the general structure of symmetry operators relevant for fusion.
In Sec. IV.2, we specialize to the simplest nontrivial case by choosing the gauge group in the action (14) to be $G=(\mathbb{Z}_{2})^{3}$. This choice allows us to explicitly demonstrate representative examples of fusion rules and to highlight characteristic features of non-invertible symmetries in a concrete and transparent setting. Guided by these examples, we present a general expression for fusion rules in Sec. IV.3, together with a practical prescription for computing fusion outcomes in a generic setup. Finally, in Sec. IV.4, we provide additional examples of fusion rules for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating how the fusion structure depends on the underlying gauge data.
@@ IV.1 Generating new conserved currents by linear combinations
from which we may identify
@@
$L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{m_{1}m_{2}m_{3}}\mathbb{P}_{m_{1}m_{2}m_{3}},$ (75)
@@ IV.2 A simple example of fusion rules of symmetry operators: $G=\left(\mathbb{Z}_{2}\right)^{3}$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]e^{{\rm i}S_{aab}}e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O},$ (78)
@@
$=$ $\frac{1}{\mathcal{Z}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=0\right]e^{{\rm i}S_{aab}}\mathcal{O}$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=w\right]e^{{\rm i}S_{aab}}\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}.$ (84)
@@
Here, $U_{m_{1}=1}L_{m_{1}=1,m_{2}=1}$ denotes the symmetry operator obtained by fusing $U_{m_{1}=1}$ and $L_{m_{1}=1,m_{2}=1}$, whose expression is $e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}+\ln\mathcal{U}_{m_{1}=1,m_{2}=1}}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$.

### v3

@@ III.4 Type-II conserved currents and non-invertible symmetry operators
Besides the type-I equations of motion discussed above, the theory also admits another class of EoMs. As an example, consider a local variation of $a^{1}$, $a^{1}\rightarrow a^{1}+\delta a^{1}$. The variation of the action is
$\delta S\left[\delta a^{1}\right]=\int_{M^{4}}\left(\frac{N_{1}}{2\pi}\,\delta a^{1}\,db^{1}+q\,\delta a^{1}a^{2}b^{3}\right).$
Requiring $\delta S\left[\delta a^{1}\right]=0\mod{2\pi}$ yields
@@
A natural question is whether one can still extract a continuity equation from the type-II EoM (26). The answer is affirmative, but only under additional conditions. The key step is to rewrite the EoM in the form of a vanishing total derivative. To this end, we introduce two auxiliary fields $\phi^{2}$ and $\beta^{3}$ such that $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. We note that $d\phi^{2}$ ($d\beta^{3}$) is an integer multiple of $a^{2}$ ($b^{3}$). Therefore, the fields $\phi^{2}$ and $\beta^{3}$ satisfy the Dirac quantization conditions $\oint d\phi^{2}\in 2\pi\mathbb{Z}$ and $\oint d\beta^{3}\in 2\pi\mathbb{Z}$ when $\oint a^{2}$ and $\oint b^{3}$ take the corresponding allowed values. With $\phi^{2}$ and $\beta^{3}$, the EoM (26) can be rewritten as
@@
Here, the two coefficients $\frac{1}{2}$ are not fixed by gauge invariance: the above equation still holds if one replaces them by two coefficients $c_{1}$ and $c_{2}$ with $c_{1}+c_{2}=1$. Rather, we split the term $\frac{2\pi q}{N_{1}}a^{2}b^{3}$ symmetrically in order to reflect that the twisted term in the action (14) is symmetric under exchanging $a^{1}$ and $a^{2}$, up to the overall antisymmetry of the wedge product. With this convention, the currents extracted from the EoMs take a symmetric form; see Table 1. Using $\frac{N_{2}}{2\pi}da^{2}=0$ and $\frac{N_{3}}{2\pi}db^{3}=0$, we further obtain
@@
$d\left(*J_{m}^{1}\right)=0,\qquad J_{m}^{1}=*\left(b^{1}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}\right),$ (29)
@@
However, an important subtlety arises: while the EoM (26) holds identically, the continuity equation (29)—and thus the conservation of $J^{1}_{m}$—is valid only when $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$ and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$ are satisfied. Therefore, for $Q_{m}^{1}$ to be conserved, these relations must hold at least on the submanifold $\sigma$. Equivalently, these conditions can be expressed as $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$, because on the closed manifold $\sigma$ one has $\int_{\sigma}d\beta^{3}=0$, and $\int_{\gamma}d\phi^{2}=0$ for any closed curve $\gamma\subset\sigma$. In other words, although $\mathcal{U}_{1}$ is a unitary operator built from $Q_{m}^{1}$, it defines a symmetry operation only when $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$ are satisfied. Accordingly, we define the symmetry operator as
@@
We further note that $L^{1}\left(\sigma\right)$ closely resembles the Wilson operator describing a loop excitation carrying one unit of $\mathbb{Z}_{N_{1}}$ gauge flux [74], up to an overall normalization factor. With this Wilson-operator picture in mind, the symmetry operator admits the following interpretation. Inserting such an operator creates a pair of loop excitations, transports one of them along $\sigma$, and finally annihilates the pair. During this process, the operator measures how many $\mathbb{Z}_{N_{1}}$ gauge charges (carried by particle excitations) are enclosed in the volume bounded by $\sigma$. This is analogous to the 22d toric code, where moving an $m$ excitation along a closed path measures the number of enclosed $e$ excitations. The enclosed number of $\mathbb{Z}_{N_{1}}$ gauge charges, $n_{e_{1}}$, is reflected by the phase factor $\exp\left({\rm i}\frac{2\pi}{N_{1}}\,n_{e_{1}}\right)$. Since the loop carries one unit of gauge flux, the measurement outcome is defined modulo $N_{1}$. The “symmetry” here is reflected in the fact that $\sigma$ can be smoothly deformed without changing $Q_{m}^{1}$, but only within the subspace where $\int_{\sigma}\frac{pN_{3}}{N_{123}}b^{3}=0$ and $\int_{\gamma\subset\sigma}\frac{pN_{2}}{N_{123}}a^{2}=0$. Outside this subspace, the symmetry need not exist. In this way, we establish a direct connection between a non-Abelian loop excitation and a non-invertible symmetry.
@@
$\mathcal{U}_{m_{1}}$ $=e^{{\rm i}Q^{1}_{m_{1}}}$
$=\exp\left[{\rm i}\left(\int_{\sigma}m_{1}b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta_{m_{1}}^{3}+\frac{1}{2}\phi_{m_{1}}^{2}\frac{N_{3}}{2\pi}b^{3}\right)\right].$ (36)
@@
In the same manner, we can obtain the EoM of the gauge field $b^{2}$,
@@
$d\left(*J_{m}^{2}\right)=0,\qquad J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right),$ (38)
@@
$L_{m_{2}}^{2}\left(\sigma\right)=\mathcal{U}_{m_{2}}\,\delta\left(\int_{\sigma}\frac{m_{2}pN_{3}}{N_{123}}b^{3}\right)\delta\left(\int_{\gamma\subset\sigma}\frac{m_{2}pN_{1}}{N_{123}}a^{1}\right),$ (41)
@@
$\mathcal{U}_{m_{2}}$ $=e^{{\rm i}\int_{\sigma}*J_{m_{2}}^{2}}$
$=\exp\left[{\rm i}\left(\int_{\sigma}m_{2}b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta_{m_{2}}^{3}-\frac{1}{2}\phi_{m_{2}}^{1}\frac{N_{3}}{2\pi}b^{3}\right)\right].$ (42)
@@
The above discussion also applies to the EoM of the gauge field $a^{3}$,
@@
$d\left(*J_{m}^{3}\right)=0,\qquad J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\phi^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right),$ (44)
@@
$\mathcal{U}_{m_{3}}$ $=e^{{\rm i}\int_{\gamma}*J_{m_{3}}^{3}}$
$=\exp\left[{\rm i}\left(\int_{\gamma}m_{3}a^{3}+\frac{1}{2}\phi_{m_{3}}^{1}\frac{N_{2}}{2\pi}a^{2}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi_{m_{3}}^{2}\right)\right].$ (48)
@@ IV Fusion rules of symmetry operators
In this section, we investigate the fusion rules of the symmetry operators identified in Sec. III. Fusion rules characterize the result of composing two symmetry operators supported on the same submanifold, and they encode the algebraic structure underlying a generalized symmetry. From a field-theoretical perspective, fusion is realized by inserting the corresponding topological operators in close succession in a correlation function.
For invertible symmetries, fusion rules are group-like: each symmetry operator is labeled by a group element, and their fusion follows the group multiplication law. By contrast, for non-invertible symmetries, fusion is intrinsically non-group-like. In particular, the product of two non-invertible symmetry operators can decompose into a linear combination of symmetry operators, rather than closing onto a single one.
To analyze fusion rules systematically, it is crucial to first identify the complete set of symmetry operators in the theory. In Sec. IV.1, we show that additional symmetry operators can be generated from new conserved currents, which themselves arise from linear combinations of existing conserved currents. This construction clarifies the general structure of symmetry operators that can appear in fusion.
In Sec. IV.2, we specialize to the simplest nontrivial case by choosing the gauge group in the action (14) to be $G=(\mathbb{Z}_{2})^{3}$. This choice allows us to present representative fusion processes explicitly and to highlight characteristic features of non-invertible symmetries in a concrete and transparent setting. Guided by these examples, we present a general expression for fusion rules in Sec. IV.3, together with a practical prescription for computing fusion outcomes in a generic setup. Finally, in Sec. IV.4, we provide additional examples for different choices of gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$, illustrating how the fusion structure depends on the underlying gauge data.
@@ IV.1 Generating new conserved currents by linear combinations
from which we may identify an conserved current
@@
$L_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)=\mathcal{U}_{m_{1}m_{2}m_{3}}\left(\sigma,\gamma_{p}\right)\mathbb{P}_{m_{1}m_{2}m_{3}},$ (75)
@@ IV.2 A simple example of fusion rules of symmetry operators: $G=\left(\mathbb{Z}_{2}\right)^{3}$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\gamma_{p}}a^{1}=w\right]e^{{\rm i}S_{\text{BR}}}e^{{\rm i}2\int_{\gamma_{p}}a^{1}}\mathcal{O},$ (78)
@@
$=$ $\frac{1}{\mathcal{Z}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=0\right]e^{{\rm i}S_{\text{BR}}}\mathcal{O}$
$=$ $\frac{1}{\mathcal{Z}}\sum_{w\in\left\{0,\pi\right\}}\int_{\text{ }}\mathcal{D}\left[a^{i},b^{i};\int_{\sigma}b^{3}=w\right]e^{{\rm i}S_{\text{BR}}}\frac{1}{2}\left(1+e^{{\rm i}\int_{\sigma}b^{3}}\right)\mathcal{O}.$ (84)
@@
Here, $U_{m_{1}=1}L_{m_{1}=1,m_{2}=1}$ denotes the symmetry operator obtained by fusing $U_{m_{1}=1}$ and $L_{m_{1}=1,m_{2}=1}$, whose expression is $e^{{\rm i}\int_{\gamma\subset\sigma}a^{1}}\mathcal{U}_{m_{1}=1,m_{2}=1}\delta\left(\int_{\gamma\subset\sigma}a^{2}-a^{1}\right)\delta\left(\int_{\sigma}b^{3}\right)$.